# Node Duplication Improves Cold-start Link Prediction

### v1

@@ Appendix D Further Experimental Results
D.2 MRR results compared with the base GNN model and cold-start baselines
Table 7 presents the performance of our methods, evaluated using MRR, compared against the base GNN model and cold-start baselines. We can observe that NodeDup(L) consistently achieves significant improvements over the baseline methods for Isolated and Low-degree nodes across all datasets. NodeDup also outperforms baselines in 13 out of 14 cases on the cold nodes. This further demonstrates the superior effectiveness of our methods in addressing cold node scenarios. Furthermore, we can also observe that our methods consistently perform on par or above baseline methods in Warm nodes and the overall setting. This observation further supports the effectiveness of our methods in maintaining and even improving the performance of Warm nodes.
Table 7: Performance compared with base GNN model and baselines for cold-start methods (evaluated by MRR). The best result is bold, and the runner-up is underlined. NodeDup and NodeDup(L) outperform GSage and cold-start baselines almost all the cases.
GSage Imbalance TailGNN Cold-brew NodeDup(L) NodeDup
Cora Isolated 16.73$\pm$1.50 17.12$\pm$0.77 20.88$\pm$0.97 15.96$\pm$1.60 22.83$\pm$0.48 25.61$\pm$1.41
Low-degree 38.46$\pm$0.62 37.93$\pm$1.17 40.19$\pm$0.96 35.20$\pm$0.55 40.20$\pm$1.02 39.78$\pm$0.97
Warm 36.97$\pm$0.60 34.94$\pm$0.87 36.39$\pm$0.51 31.97$\pm$0.31 36.99$\pm$0.41 35.34$\pm$0.32
Overall 35.91$\pm$0.51 34.59$\pm$0.81 36.49$\pm$0.59 31.84$\pm$0.17 36.89$\pm$0.47 35.82$\pm$0.34
Citeseer Isolated 29.36$\pm$2.30 28.35$\pm$1.02 22.49$\pm$1.67 21.91$\pm$5.24 34.19$\pm$0.77 38.26$\pm$1.26
Low-degree 44.13$\pm$0.38 44.67$\pm$0.44 43.92$\pm$1.55 34.65$\pm$10.10 51.58$\pm$0.56 53.71$\pm$0.64
Warm 46.68$\pm$0.48 46.95$\pm$1.01 45.93$\pm$1.17 36.45$\pm$7.50 48.89$\pm$0.53 48.05$\pm$0.42
Overall 42.60$\pm$0.59 42.72$\pm$0.52 40.87$\pm$1.34 33.13$\pm$7.90 47.00$\pm$0.44 48.05$\pm$0.54
CS Isolated 35.54$\pm$0.74 29.61$\pm$1.62 30.32$\pm$0.92 32.35$\pm$0.77 42.22$\pm$1.41 44.94$\pm$0.60
Low-degree 56.18$\pm$0.81 57.44$\pm$0.68 46.66$\pm$0.61 42.67$\pm$0.26 61.20$\pm$0.64 61.65$\pm$0.84
Warm 58.18$\pm$0.84 57.03$\pm$0.77 48.32$\pm$0.44 43.71$\pm$0.41 59.94$\pm$0.54 58.67$\pm$0.72
Overall 57.73$\pm$0.83 56.72$\pm$0.73 47.96$\pm$0.45 43.48$\pm$0.38 59.83$\pm$0.52 58.74$\pm$0.70
Physics Isolated 27.73$\pm$1.10 33.61$\pm$0.34 23.17$\pm$0.74 30.62$\pm$0.30 41.12$\pm$1.10 45.62$\pm$2.45
Low-degree 61.40$\pm$0.52 62.74$\pm$0.27 47.05$\pm$0.39 41.95$\pm$0.15 64.04$\pm$0.43 65.94$\pm$0.21
Warm 66.72$\pm$0.47 66.03$\pm$0.09 55.77$\pm$0.49 46.06$\pm$0.12 66.94$\pm$0.49 66.83$\pm$0.04
Overall 66.39$\pm$0.47 65.80$\pm$0.09 55.36$\pm$0.49 45.86$\pm$0.12 66.74$\pm$0.49 66.72$\pm$0.04
Computers Isolated 4.50$\pm$0.75 5.01$\pm$0.71 4.88$\pm$0.54 4.07$\pm$0.46 8.59$\pm$1.45 9.65$\pm$1.10
Low-degree 26.65$\pm$0.62 26.85$\pm$0.31 21.22$\pm$0.56 23.40$\pm$0.59 28.85$\pm$1.13 29.78$\pm$0.32
Warm 34.11$\pm$0.40 33.77$\pm$0.17 31.02$\pm$0.34 28.75$\pm$0.23 35.11$\pm$0.31 35.63$\pm$0.14
Overall 33.98$\pm$0.40 33.65$\pm$0.16 30.88$\pm$0.34 28.64$\pm$0.23 35.00$\pm$0.31 35.52$\pm$0.13
Photos Isolated 3.99$\pm$0.52 4.79$\pm$1.38 5.78$\pm$0.94 6.49$\pm$0.98 8.23$\pm$1.10 7.90$\pm$1.55
Low-degree 25.10$\pm$1.35 24.60$\pm$1.20 20.41$\pm$1.29 21.54$\pm$0.35 27.90$\pm$0.90 26.90$\pm$1.29
Warm 34.90$\pm$0.57 33.03$\pm$0.47 30.79$\pm$0.63 29.40$\pm$0.23 36.84$\pm$0.55 35.69$\pm$0.43
Overall 34.71$\pm$0.57 32.87$\pm$0.47 30.60$\pm$0.63 29.26$\pm$0.22 36.66$\pm$0.54 35.52$\pm$0.43
IGB-100K Isolated 53.20$\pm$0.24 50.81$\pm$0.41 45.25$\pm$0.26 48.42$\pm$0.25 59.34$\pm$0.51 61.75$\pm$0.47
Low-degree 55.93$\pm$0.28 55.79$\pm$0.30 51.11$\pm$0.29 51.92$\pm$0.15 62.35$\pm$0.49 63.91$\pm$0.26
Warm 61.31$\pm$0.49 60.63$\pm$0.40 55.91$\pm$0.18 50.88$\pm$0.20 61.56$\pm$0.48 61.24$\pm$0.19
Overall 60.05$\pm$0.43 59.40$\pm$0.36 54.65$\pm$0.20 50.97$\pm$0.17 61.61$\pm$0.48 61.73$\pm$0.21
D.3 Additional results compared with augmentation baselines
Figure 7 presents the performance compared with augmentation methods on the remaining datasets. On Cora and CS datasets, we can consistently observe that NodeDup and NodeDup(L) outperform all the graph augmentation baselines for Isolated and Low-degree nodes. Moreover, for Warm nodes, NodeDup can also perform on par or above baselines. On the Computers and Photos datasets, our methods generally achieve comparable or superior performance compared to the baselines, except in comparison to TuneUP. However, it is worth noting that both NodeDup and NodeDup(L) exhibit more than 2$\times$ faster execution speed than TuneUP on these two datasets.
@@ D.3 Additional results compared with augmentation baselines
Figure 7: Performance and time-consuming compared with augmentation methods (Remaining results of Figure 5). The left histograms show the performance results, and the right histograms show the preprocessing and training time consumption of each method.
D.4 Additional results under the inductive setting
@@
We further evaluate and present the effectiveness of our methods under the inductive setting on the remaining datasets in Table 8. We can observe that both NodeDup and NodeDup(L) consistently outperform GSage for Isolated, Low-degree, and Warm nodes. Compared to NodeDup(L), NodeDup is particularly beneficial for this inductive setting.
@@
Table 8: Performance in inductive settings (Remaining results of Table 2). The best result is bold, and the runner-up is underlined. Our methods consistently outperform GSage.
@@
D.5 Performance compared with Heuristic Methods
We compare our method with traditional link prediction baselines, such as common neighbors (CN), Adamic-Adar(AA), Resource allocation (RA). The results are shown in Table 9. We observe that NodeDup can consistently outperform these heuristic methods across all the datasets, with particularly significant improvements observed on Isolated nodes.
Table 9: Performance compared with heuristic methods and DegFairGNN (Liu et al., 2023). The best result is bold. NodeDup, consistently outperforms all the heuristic methods and DegFairGNN.
CN AA RA DegFairGNN GSage NodeDup
Cora Isolated 0.00 0.00 0.00 18.70$\pm$1.53 32.20$\pm$3.58 44.27$\pm$3.82
Low-degree 20.30 20.14 20.14 38.43$\pm$0.14 59.45$\pm$1.09 61.98$\pm$1.14
Warm 38.33 38.90 38.90 42.49$\pm$1.82 61.14$\pm$0.78 59.07$\pm$0.68
Overall 25.27 25.49 25.49 39.24$\pm$1.10 58.31$\pm$0.68 58.92$\pm$0.82
Citeseer Isolated 0.00 0.00 0.00 15.50$\pm$1.27 47.13$\pm$2.43 57.54$\pm$1.04
Low-degree 26.86 27.00 27.00 45.06$\pm$0.96 61.88$\pm$0.79 75.50$\pm$0.39
Warm 37.30 39.02 39.02 55.47$\pm$1.08 71.45$\pm$0.52 74.68$\pm$0.67
Overall 30.81 31.85 31.85 44.58$\pm$1.03 63.77$\pm$0.83 71.73$\pm$0.47
CS Isolated 0.00 0.00 0.00 17.93$\pm$1.35 56.41$\pm$1.61 65.87$\pm$1.70
Low-degree 39.60 39.60 39.60 49.83$\pm$0.68 75.95$\pm$0.25 81.12$\pm$0.36
Warm 72.73 72.74 72.72 61.72$\pm$0.37 84.37$\pm$0.46 84.76$\pm$0.41
Overall 69.10 69.11 69.10 60.20$\pm$0.37 83.33$\pm$0.42 84.23$\pm$0.39
Physics Isolated 0.00 0.00 0.00 19.48$\pm$2.94 47.41$\pm$1.38 66.65$\pm$0.95
Low-degree 46.08 46.08 46.08 47.63$\pm$0.52 79.31$\pm$0.28 84.04$\pm$0.22
Warm 85.48 85.74 85.70 62.79$\pm$0.82 90.28$\pm$0.23 90.33$\pm$0.05
Overall 83.87 84.12 84.09 62.13$\pm$0.76 89.76$\pm$0.22 90.03$\pm$0.05
Computers Isolated 0.00 0.00 0.00 9.36$\pm$1.81 9.32$\pm$1.44 19.62$\pm$2.63
Low-degree 28.31 28.31 28.31 18.90$\pm$0.81 57.91$\pm$0.97 61.16$\pm$0.92
Warm 59.67 63.50 62.84 31.44$\pm$2.25 66.87$\pm$0.47 68.10$\pm$0.25
Overall 59.24 63.03 62.37 31.27$\pm$2.22 66.67$\pm$0.47 67.94$\pm$0.25
Photos Isolated 0.00 0.00 0.00 12.99$\pm$1.51 9.25$\pm$2.31 17.84$\pm$3.53
Low-degree 28.44 28.78 28.78 20.18$\pm$0.21 52.61$\pm$0.88 54.13$\pm$1.58
Warm 64.53 67.26 66.88 42.72$\pm$0.89 67.64$\pm$0.55 68.68$\pm$0.49
Overall 63.94 66.64 66.26 42.37$\pm$0.87 67.32$\pm$0.54 68.39$\pm$0.48
IGB-100K Isolated 0.00 0.00 0.00 57.09$\pm$21.08 75.92$\pm$0.52 88.04$\pm$0.20
Low-degree 12.26 12.26 12.26 59.45$\pm$21.84 79.38$\pm$0.23 88.98$\pm$0.17
Warm 30.65 30.65 30.65 65.57$\pm$20.43 86.42$\pm$0.24 88.28$\pm$0.20
Overall 26.22 26.22 26.22 64.16$\pm$20.70 84.77$\pm$0.21 88.39$\pm$0.18
D.6 Performance compared with the methods tackling degree bias in GNNs
We have conducted the experiments to compare with the existing methods, TailGNN (Liu et al., 2021) and Cold-brew (Zheng et al., 2021), which also tackle degree bias in GNNs. The results, presented in Table 1, consistently show our methods outperforming these approaches. DegFairGNN (Liu et al., 2023) introduces a learnable debiasing function in the GNN architecture to produce fair representations for nodes, aiming for similar predictions for nodes within the same class, regardless of their degrees. Unfortunately, we’ve found that this approach is not well-suited for link prediction tasks for several reasons: (1) This method is designed specifically for node classification tasks. For example, the fairness loss, which ensures prediction distribution uniformity among low and high-degree node groups, is not suitable for link prediction because there is no defined node class in link prediction tasks. (2) This approach achieves significant performance in node classification tasks by effectively mitigating degree bias. However, in the context of link prediction, the degree trait is crucial. Applying DegFairGNN (Liu et al., 2023) would compromise the model’s ability to learn from structural information, such as isomorphism and common neighbors. This, in turn, would negatively impact link prediction performance, as evidenced by references (Zhang & Chen, 2018; Chamberlain et al., 2022).
D.7 Performance compared with upsampling
In Section 3, we discussed the issue of under-representation of cold nodes during the training of LP, which is the main cause of their unsatisfactory performance. To tackle this problem, one straightforward and naive approach is upsampling (Provost, 2000), which involves increasing the number of samples in the minority class. In order to further demonstrate the effectiveness of our methods, we conducted experiments where we doubled the edge sampling probability of code nodes, aiming to enhance their visibility. The results are presented in Table 10. We can observe that NodeDup(L) consistently outperforms upsampling, and NodeDup outperforms upsampling in almost all the cases, except for Warm nodes on Cora.
Table 10: Performance compared with upsampling(evaluated by Hits@10). The best result is bold, and the runner-up is underlined. NodeDup(L) consistently outperforms upsampling.
Dataset Method Isolated Low-degree Warm Overall
Cora Upsampling 32.81$\pm$2.75 59.57$\pm$0.60 60.49$\pm$0.81 57.90$\pm$0.65
NodeDup(L) 39.76$\pm$1.32 62.53$\pm$1.03 62.07$\pm$0.37 60.49$\pm$0.49
NodeDup 44.27$\pm$3.82 61.98$\pm$1.14 59.07$\pm$0.68 58.92$\pm$0.82
Citeseer Upsampling 46.88$\pm$0.45 62.32$\pm$1.57 71.33$\pm$1.35 63.81$\pm$0.81
NodeDup(L) 52.46$\pm$1.16 73.71$\pm$1.22 74.99$\pm$0.37 70.34$\pm$0.35
NodeDup 57.54$\pm$1.04 75.50$\pm$0.39 74.68$\pm$0.67 71.73$\pm$0.47
CS Upsampling 49.63$\pm$2.24 75.62$\pm$0.13 83.40$\pm$0.73 82.34$\pm$0.64
NodeDup(L) 65.18$\pm$1.25 81.46$\pm$0.57 85.48$\pm$0.26 84.90$\pm$0.29
NodeDup 65.87$\pm$1.70 81.12$\pm$0.36 84.76$\pm$0.41 84.23$\pm$0.39
Physics Upsampling 52.01$\pm$0.97 79.63$\pm$0.13 89.41$\pm$0.32 89.33$\pm$0.46
NodeDup(L) 65.04$\pm$0.63 82.70$\pm$0.22 90.44$\pm$0.23 90.09$\pm$0.22
NodeDup 66.65$\pm$0.95 84.04$\pm$0.22 90.33$\pm$0.05 90.03$\pm$0.05
Computers Upsampling 11.36$\pm$0.72 58.23$\pm$0.88 67.07$\pm$0.49 66.87$\pm$0.48
NodeDup(L) 17.11$\pm$1.62 62.14$\pm$1.06 68.02$\pm$0.41 67.86$\pm$0.41
NodeDup 19.62$\pm$2.63 61.16$\pm$0.92 68.10$\pm$0.25 67.94$\pm$0.25
Photos Upsampling 10.92$\pm$2.15 51.67$\pm$0.98 65.75$\pm$0.73 65.45$\pm$0.71
NodeDup(L) 21.50$\pm$2.14 55.70$\pm$1.38 69.68$\pm$0.87 69.40$\pm$0.86
NodeDup 17.84$\pm$3.53 54.13$\pm$1.58 68.68$\pm$0.49 68.39$\pm$0.48
IGB-100K Upsampling 75.49$\pm$0.90 79.47$\pm$0.11 86.54$\pm$0.19 84.87$\pm$0.14
NodeDup(L) 87.43$\pm$0.44 88.37$\pm$0.24 88.54$\pm$0.31 88.47$\pm$0.28
NodeDup 88.04$\pm$0.20 88.98$\pm$0.17 88.28$\pm$0.20 88.39$\pm$0.18
Table 11: Performance on the large-scale dataset. The best result is bold. Our method consistently outperforms GSage on IGB-1M.
GSage NodeDup
IGB-1M Isolated 82.10$\pm$0.06 87.81$\pm$0.40
Low-degree 84.73$\pm$0.06 90.84$\pm$0.03
Warm 89.98$\pm$0.02 91.31$\pm$0.02
Overall 89.80$\pm$0.02 91.29$\pm$0.02
D.8 Performance on Large-scale datasets
As outlined in Section 3.1, our methods incur a minimal increase in time complexity compared to base GNNs, with the increase being linearly proportional to the number of cold nodes. This ensures scalability. Besides, the effectiveness of our method is also insensitive to dataset size. We extend our experiments to the IGB-1M dataset, featuring 1 million nodes and 12 million edges. The findings, which we detail in Table 11, affirm the effectiveness of our methods in handling large-scale datasets, consistent with observations from smaller datasets.
Table 12: Performance with different encoders (Remaining results of Table 3), where the inner product is the decoder. The best result for each encoder is bold, and the runner-up is underlined. Our methods consistently outperform the base models, particularly for Isolated and Low-degree nodes.
@@ D.8 Performance on Large-scale datasets
Table 13: Link prediction performance with MLP decoder (Remaining results of Table 4), where GSage is the encoder. Our methods achieve better performance than the base model.

### v2

@@ Appendix C Further Experimental Results
C.2 Further analysis about the duplication node types
To make our analysis about the duplication node types more comprehensive, we conducted additional experiments to evaluate different node duplication strategies under both the NodeDup and NodeDup(L) settings. We present detailed experiments on the Citeseer dataset, where we duplicate different subsets of nodes: isolated nodes, cold nodes, mid-warm nodes, warm nodes, random nodes, and all nodes. We report the corresponding training time, memory usage, and performance across multiple evaluation subsets. The results are summarized in Table˜8.
From the table, we can observe that for both NodeDup and NodeDup(L), duplicating mid-warm and warm nodes results in little to no performance improvement. Random duplication provides clear improvements over both no duplication and warm node duplication, but remains less effective than cold node duplication. Since random duplication and cold nodes duplication duplicate the same number of nodes, they incur similar memory and training costs. Under the NodeDup setting, compared to cold nodes duplication, all nodes duplication increases training time by approximately 32% (3.3s vs. 2.5s) and memory usage by approximately 35% (93.99MB vs. 69.46MB), while achieving only marginal additional performance gain (Low-degree Hits@10: 76.09 vs. 75.50). Under the NodeDup(L) setting, full duplication increases training time by approximately 14% (2.4s vs. 2.1s), while memory usage remains identical, again with limited additional performance gain (Low-degree Hits@10: 75.32 vs. 74.99).
These results demonstrate that our proposed selective cold node duplication achieves most of the performance improvement while significantly reducing both training time and memory cost compared to full duplication. This further supports the practical advantage of our method under constrained compute budgets.
Table 8: Computation time, memory usage, and link prediction performance with different duplication nodes of NodeDup and NodeDup(L) on Citeseer. "D_*" indicates duplication of "*" group nodes for one time.
NodeDup
Time(s) Memory(MB) Performance (Hits@10)
Training Node attributes Edges Isolated Low-degree Warm Overall
Supervised 1.8 47.00 0.14 47.13 61.88 71.45 63.77
D_Isolated 1.9 54.03 0.14 54.04 72.28 74.53 69.95
D_Cold 2.4 69.46 0.14 57.54 75.50 74.68 71.73
D_Mid-warm 2.2 55.83 0.14 46.93 61.34 71.84 63.75
D_Warm 2.2 57.53 0.14 47.49 62.20 71.54 63.99
D_Random 2.4 69.46 0.14 54.10 72.39 75.05 70.06
D_All 3.3 93.99 0.14 58.87 76.09 76.01 72.44
NodeDup(L)
Time(s) Memory(MB) Performance(Hits@10)
Training Node attributes Edges Isolated Low-degree Warm Overall
Supervised 1.8 47.00 0.14 47.13 61.88 71.45 63.77
D_Isolated 2.0 47.00 0.14 49.06 69.95 74.67 68.32
D_Cold 2.1 47.00 0.14 52.46 73.71 74.99 70.34
D_Mid-warm 2.2 47.00 0.14 46.05 61.57 72.30 63.88
D_Warm 2.1 47.00 0.14 46.53 61.86 71.77 63.82
D_Random 2.2 47.00 0.14 48.92 69.59 73.90 67.81
D_All 2.4 47.00 0.14 52.44 74.05 75.21 70.49
Table 9: Performance on the large-scale dataset. The best result is bold. Our method consistently outperforms GSage on IGB1M.
GSage NodeDup
IGB1M Isolated 82.10$\pm$0.06 87.81$\pm$0.40
Low-degree 84.73$\pm$0.06 90.84$\pm$0.03
Warm 89.98$\pm$0.02 91.31$\pm$0.02
Overall 89.80$\pm$0.02 91.29$\pm$0.02
C.3 Performance on large-scale datasets
As outlined in Section˜3.1, our methods incur a minimal increase in time complexity compared to base GNNs, with the increase being linearly proportional to the number of cold nodes. This ensures scalability. Besides, the effectiveness of our method is also insensitive to dataset size. We extend our experiments to the IGB1M dataset, featuring 1 million nodes and 12 million edges. The findings, which we detail in Table˜9, affirm the effectiveness of our methods in handling large-scale datasets, consistent with observations from smaller datasets.
Table 10: Performance on heterophilic datasets. The best result for each dataset is bold.
GSage NodeDup(L) NodeDup
Chameleon Isolated 24.91$\pm$6.75 30.76$\pm$4.02 27.37$\pm$2.88
Low-degree 79.09$\pm$1.21 80.11$\pm$0.68 80.91$\pm$0.41
Warm 94.00$\pm$0.23 94.01$\pm$0.12 93.68$\pm$0.44
Overall 92.77$\pm$0.19 92.88$\pm$0.10 92.57$\pm$0.44
Squirrel Isolated 25.05$\pm$3.70 33.07$\pm$3.20 30.11$\pm$1.57
Low-degree 63.34$\pm$2.12 66.61$\pm$0.26 68.05$\pm$0.80
Warm 93.35$\pm$0.22 93.43$\pm$0.11 93.82$\pm$0.13
Overall 92.89$\pm$0.23 93.02$\pm$0.11 93.41$\pm$0.13
C.4 Performance on heterophily datasets
We have conducted experiments on two heterophilic datasets (i.e., Chameleon (Pei et al., 2020) and Squirrel (Pei et al., 2020)), with the results shown in Table˜10. Our methods improve GNN performance across all settings on these datasets. Specifically, NodeDup and NodeDup(L) enhance the performance of Isolated nodes by 9.9% and 23.5% on Chameleon, and by 20.2% and 32.0% on Squirrel.
Table 11: Performance on the MovieLens. The best result is bold. Our method consistently outperforms GSage.
MovieLens_1M
GSage NodeDup(L) NodeDup
Isolated 0 3.08 5.38
Low_degree 30.7 35.07 37.69
Warm 41.79 44.64 45.52
Overall 41.71 44.56 45.78
C.5 Performance on recommendation datasets
To further evaluate the practical applicability of our method to real-world recommendation scenarios, we followed the approach introduced in KGCN (Huang et al., 2021) to construct a movie-movie graph, where two movies are connected if they received high ratings from the same user. This graph is then used to recommend similar movies to users. This setup forms an item-based collaborative filtering recommendation task, allowing us to apply our methods. The results is shown in Table˜11. Compared to the baseline, both NodeDup and NodeDup(L) achieve consistent improvements, particularly on low-degree and isolated nodes, where these cold-start items often limits recommendation accuracy.
C.6 Performance compared with heuristic methods
We compare our method with traditional link prediction baselines, such as common neighbors (CN), Adamic-Adar(AA), Resource allocation (RA). The results are shown in Table˜12. We observe that NodeDup can consistently outperform these heuristic methods across all the datasets, with particularly significant improvements observed on Isolated nodes.
Table 12: Performance compared with heuristic methods, Upsampling and DegFairGNN (Liu et al., 2023). The best result is bold.
CN AA RA Upsampling DegFairGNN GSage NodeDup
Cora Isolated 0.00 0.00 0.00 32.81$\pm$2.75 18.70$\pm$1.53 32.20$\pm$3.58 44.27$\pm$3.82
Low-degree 20.30 20.14 20.14 59.57$\pm$0.60 38.43$\pm$0.14 59.45$\pm$1.09 61.98$\pm$1.14
Warm 38.33 38.90 38.90 60.49$\pm$0.81 42.49$\pm$1.82 61.14$\pm$0.78 59.07$\pm$0.68
Overall 25.27 25.49 25.49 57.90$\pm$0.65 39.24$\pm$1.10 58.31$\pm$0.68 58.92$\pm$0.82
Citeseer Isolated 0.00 0.00 0.00 46.88$\pm$0.45 15.50$\pm$1.27 47.13$\pm$2.43 57.54$\pm$1.04
Low-degree 26.86 27.00 27.00 62.32$\pm$1.57 45.06$\pm$0.96 61.88$\pm$0.79 75.50$\pm$0.39
Warm 37.30 39.02 39.02 71.33$\pm$1.35 55.47$\pm$1.08 71.45$\pm$0.52 74.68$\pm$0.67
Overall 30.81 31.85 31.85 63.81$\pm$0.81 44.58$\pm$1.03 63.77$\pm$0.83 71.73$\pm$0.47
CS Isolated 0.00 0.00 0.00 49.63$\pm$2.24 17.93$\pm$1.35 56.41$\pm$1.61 65.87$\pm$1.70
Low-degree 39.60 39.60 39.60 75.62$\pm$0.13 49.83$\pm$0.68 75.95$\pm$0.25 81.12$\pm$0.36
Warm 72.73 72.74 72.72 83.40$\pm$0.73 61.72$\pm$0.37 84.37$\pm$0.46 84.76$\pm$0.41
Overall 69.10 69.11 69.10 82.34$\pm$0.64 60.20$\pm$0.37 83.33$\pm$0.42 84.23$\pm$0.39
Physics Isolated 0.00 0.00 0.00 52.01$\pm$0.97 19.48$\pm$2.94 47.41$\pm$1.38 66.65$\pm$0.95
Low-degree 46.08 46.08 46.08 79.63$\pm$0.13 47.63$\pm$0.52 79.31$\pm$0.28 84.04$\pm$0.22
Warm 85.48 85.74 85.70 89.41$\pm$0.32 62.79$\pm$0.82 90.28$\pm$0.23 90.33$\pm$0.05
Overall 83.87 84.12 84.09 89.33$\pm$0.46 62.13$\pm$0.76 89.76$\pm$0.22 90.03$\pm$0.05
Computers Isolated 0.00 0.00 0.00 11.36$\pm$0.72 9.36$\pm$1.81 9.32$\pm$1.44 19.62$\pm$2.63
Low-degree 28.31 28.31 28.31 58.23$\pm$0.88 18.90$\pm$0.81 57.91$\pm$0.97 61.16$\pm$0.92
Warm 59.67 63.50 62.84 67.07$\pm$0.49 31.44$\pm$2.25 66.87$\pm$0.47 68.10$\pm$0.25
Overall 59.24 63.03 62.37 66.87$\pm$0.48 31.27$\pm$2.22 66.67$\pm$0.47 67.94$\pm$0.25
Photos Isolated 0.00 0.00 0.00 10.92$\pm$2.15 12.99$\pm$1.51 9.25$\pm$2.31 17.84$\pm$3.53
Low-degree 28.44 28.78 28.78 51.67$\pm$0.98 20.18$\pm$0.21 52.61$\pm$0.88 54.13$\pm$1.58
Warm 64.53 67.26 66.88 65.75$\pm$0.73 42.72$\pm$0.89 67.64$\pm$0.55 68.68$\pm$0.49
Overall 63.94 66.64 66.26 65.45$\pm$0.71 42.37$\pm$0.87 67.32$\pm$0.54 68.39$\pm$0.48
IGB-100K Isolated 0.00 0.00 0.00 75.49$\pm$0.90 57.09$\pm$21.08 75.92$\pm$0.52 88.04$\pm$0.20
Low-degree 12.26 12.26 12.26 79.47$\pm$0.11 59.45$\pm$21.84 79.38$\pm$0.23 88.98$\pm$0.17
Warm 30.65 30.65 30.65 86.54$\pm$0.19 65.57$\pm$20.43 86.42$\pm$0.24 88.28$\pm$0.20
Overall 26.22 26.22 26.22 84.87$\pm$0.14 64.16$\pm$20.70 84.77$\pm$0.21 88.39$\pm$0.18
C.7 Performance compared with additional cold-start methods
Upsampling (Provost, 2000). In Section˜3, we discussed the issue of under-representation of cold nodes during the training of LP, which is the main cause of their unsatisfactory performance. To tackle this problem, one straightforward and naive approach is upsampling (Provost, 2000), which involves increasing the number of samples in the minority class. In order to further demonstrate the effectiveness of our methods, we conducted experiments where we doubled the edge sampling probability of code nodes, aiming to enhance their visibility. The results are presented in Table˜12. We can observe that NodeDup outperforms upsampling in almost all the cases, except for Warm nodes on Cora.
Table 13: Performance compared with GRADE (Wang et al., 2022) and SAILOR (Liao et al., 2023). The best result is bold.
GCN GRADE SAILOR NodeDup(L) NodeDup
Cora Isolated 40.61$\pm$3.52 43.29$\pm$2.62 45.12$\pm$1.29 42.93$\pm$2.68 46.71$\pm$1.53
Low-degree 63.86$\pm$0.78 58.76$\pm$1.27 62.98$\pm$3.92 64.63$\pm$1.60 64.10$\pm$1.37
Warm 60.59$\pm$0.62 60.00$\pm$0.51 57.34$\pm$3.80 61.31$\pm$0.43 60.26$\pm$0.70
Overall 60.16$\pm$0.44 56.90$\pm$0.71 58.33$\pm$3.51 61.02$\pm$0.61 59.90$\pm$0.89
Citeseer Isolated 45.56$\pm$1.30 50.11$\pm$2.24 49.29$\pm$2.75 47.84$\pm$0.94 50.64$\pm$1.10
Low-degree 69.37$\pm$0.36 59.49$\pm$1.13 65.78$\pm$1.11 70.15$\pm$1.56 71.13$\pm$0.64
Warm 74.68$\pm$0.38 70.01$\pm$0.50 72.66$\pm$0.37 73.26$\pm$0.97 72.93$\pm$0.78
Overall 67.48$\pm$0.42 61.11$\pm$0.72 64.80$\pm$0.66 67.47$\pm$0.83 67.67$\pm$0.66
The methods tackling degree bias in GNNs. SAILOR (Liao et al., 2023) proposes a structural augmentation framework to enhance the representation learning of tail nodes. GRADE (Wang et al., 2022) improves structural fairness using graph contrastive learning methods. We used GCN as the encoder for both NodeDup(L) and NodeDup to ensure consistency, as both GRADE and SAILOR used GCN as their encoder. Table˜13 shows that our methods outperform these baselines in all settings. Additionally, both GRADE and SAILOR perform better than vanilla GCN on Isolated nodes, which is the primary focus of their training. The sub-optimal performance of GRADE and SAILOR on Low-degree nodes in link prediction likely stems from the fact that their augmentation strategies, which are tailored for node classification, are less suited for ranking-based link prediction tasks. The specific reasons are as follows: GRADE encourages representation smoothness via degree-aware contrastive learning, which benefits node classification but reduces the embedding discrimination required for link prediction, particularly when ranking edges involving low-degree nodes, as reflected by Hits@10. SAILOR constructs pseudo-homophilic edges based on node labels; however, label similarity does not always align with link formation in link prediction, and adding label-based edges may introduce noise.
DegFairGNN (Liu et al., 2023) introduces a learnable debiasing function in the GNN architecture to produce fair representations for nodes, aiming for similar predictions for nodes within the same class, regardless of their degrees. Unfortunately, we’ve found in Table˜12 that this approach is not well-suited for link prediction tasks for several reasons: (1) This method is designed specifically for node classification tasks. For example, the fairness loss, which ensures prediction distribution uniformity among low and high-degree node groups, is not suitable for link prediction because there is no defined node class in link prediction tasks. (2) This approach achieves significant performance in node classification tasks by effectively mitigating degree bias. However, in the context of link prediction, the degree trait is crucial. Applying DegFairGNN (Liu et al., 2023) would compromise the model’s ability to learn from structural information, such as isomorphism and common neighbors. This, in turn, would negatively impact link prediction performance, as evidenced by references (Zhang & Chen, 2018; Chamberlain et al., 2022).
@@ C.7 Performance compared with additional cold-start methods
Figure 11: Time-consuming compared with cold-start methods. The histograms show the preprocessing and training time consumption of each method.
C.8 Efficiency comparison with the base GNN model and cold-start baselines
The efficiency comparison between our methods and cold-start baselines is presented in Figure˜11. We can observe that our methods and Imbalance exhibit similar efficiency, comparable to GSage. However, TailGNN and Cold-brew demand significantly more preprocessing and training time. Cold-brew, in particular, needs the most preprocessing time as it needs to train a teacher model for distillation.
C.9 Additional results compared with augmentation baselines
Figure˜12 presents the performance compared with augmentation methods on the remaining datasets. On Cora and CS datasets, we can consistently observe that NodeDup and NodeDup(L) outperform all the graph augmentation baselines for Isolated and Low-degree nodes. Moreover, for Warm nodes, NodeDup can also perform on par or above baselines. On the Computers and Photos datasets, our methods generally achieve comparable or superior performance compared to the baselines, except in comparison to TuneUP. However, it is worth noting that both NodeDup and NodeDup(L) exhibit more than 2$\times$ faster execution speed than TuneUP on these two datasets.
Figure 12: Performance and time-consuming compared with augmentation methods (Remaining results of Figure˜8). The left histograms show the performance results, and the right histograms show the preprocessing and training time consumption of each method.
C.10 Additional results under the inductive setting
@@
We further evaluate and present the effectiveness of our methods under the inductive setting on the remaining datasets in Table˜14. We can observe that both NodeDup and NodeDup(L) consistently outperform GSage for Isolated, Low-degree, and Warm nodes. Compared to NodeDup(L), NodeDup is particularly beneficial for this inductive setting.
@@
Table 14: Performance in inductive settings (Remaining results of Table˜2). The best result is bold, and the runner-up is underlined. Our methods consistently outperform GSage.
@@
Table 15: Performance with different encoders (Remaining results of Table˜3), where the inner product is the decoder. The best result for each encoder is bold, and the runner-up is underlined. Our methods consistently outperform the base models, particularly for Isolated and Low-degree nodes.
@@
Table 16: Link prediction performance with MLP decoder (Remaining results of Table˜4), where GSage is the encoder. Our methods achieve better performance than the base model.